Marginal cost is the change in total cost caused by a one-unit increase in output.

A) True

B) False

Average total cost is equal to average variable cost plus average fixed cost.

A) True

B) False

Diminishing marginal returns means that as you combine more units of a variable resource with a set of fixed resources, the marginal product decreases.

A) True

B) False
Answer #1

Answer

True

marginal cost is the change in total cost because of an increase in one unit of output.

----------

True

ATC=AVC+AFC

The total cost is the sum of fixed cost and variable cost. Average cost =TC/Q. and AFC=FC/Q, AVC=VC/Q

----------

True

A diminishing return is decreased in the marginal product of variable input because of the capacity of a variable input is restricted by the fixed input as both are used in combinations.